A candy mixture is created with 2 types of candy, one costing $4 per pound and the other $3.50 per pound. How much of each type is needed for a 5 pound box that costs $18?
Answer by stanbon(75887)

A candy mixture is created with 2 types of candy, one costing $4 per pound and the other $3.50 per pound. How much of each type is needed for a 5 pound box that costs $18?
------------------
Equations:
value + value = value
4x + 3.50(5-x) = 18
Multiply thru by 100 to get:
400x + 350*5 - 350x = 1800
50x = 50
x = 1 lb (amt. of $4.00 candy needed in the mixture)
5-x = 4 lb (amt. of $3.50 candy needed in the mixture)
======================
Cheers,
Stan H.